OpenStudy (ibrafacts):

Is this right I got 18 for the area

OpenStudy (ibrafacts):

#37

OpenStudy (jdoe0001):

recall that the area of a triangle is 1/2 base * height thus notice, the internal corners of the triangles in the rhombus, have 45 degrees whilst the diagonals are perpendicular by definition, meaning 90 degrees thus  so, using the 45-45-90 rule

OpenStudy (ibrafacts):

yup

OpenStudy (ibrafacts):

hyp is 3sqrt2

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

hmmm wait a sec... actually... yes 3 * 3 = 9 for a half and 3 * 3 = 9 for the second half, 9 + 9 = 18 yes

563blackghost (563blackghost):

This is a 45-45-90 degree triangle which means that the legs are both equal so since one side is 3cm the other will be as well....So we need to figure out both diameters in which we multiply 3 by 2... \(\Large{3 \times 2 = diameter1=6}\) \(\Large{3 \times 2 = diameter2=6}\) In which we input into this equation... \(\LARGE{A=\frac{1}{2}d1d2}\) Input... \(\LARGE{A=\frac{1}{2}6(6)}\)simplify... \(\LARGE{A=\frac{1}{2}36}\) in which we divide by 2... \(\Huge\color{red}{\frac{36}{2}=18}\) So yes that would be the answer Is this right @jdoe0001

OpenStudy (ibrafacts):

oh

OpenStudy (ibrafacts):

thank you

OpenStudy (jdoe0001):

@563blackghost yeap

OpenStudy (ibrafacts):

would simple radical form be 3sqrt6?

563blackghost (563blackghost):

I dont know O.o We werent trying to find the square root of 18 so I guess this is how it would be.... @jdoe0001

OpenStudy (ibrafacts):

oh but thank you still for trying:)

563blackghost (563blackghost):

np ^^